# A Regret Perspective on Online Selective Generation

### v1

@@ G.4 A Proof of Theorem 2
Also, to avoid notational clutter, we define
[…]
In the following, we find the lower and upper bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$ to combine them for the final regret bound, where the expectation is taken over $\tau_{t}\sim p_{t}$ from samples drawn in advance by an oblivious adversary.
The upper bound of potentials.
We find an upper bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$. To this end, we first consider an upper bound of $\mathbbm{E}\ln\frac{W_{t}}{W_{t-1}}$. In particular, we have
$\ln\frac{W_{t}}{W_{t-1}}$ $=\ln\frac{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t}^{i})}{\sum_{j=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{j})}$
$=\ln\frac{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{i})\,\exp(-\eta\tilde{\ell}_{t}^{i})}{\sum_{j=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{j})}$
$=\ln\sum_{i=1}^{|\mathcal{H}|}\frac{\exp(-\eta\tilde{L}_{t-1}^{i})}{\sum_{j=1}^{|\mathcal{H}|}\exp(-\eta\tilde{L}_{t-1}^{j})}\exp(-\eta\tilde{\ell}_{t}^{i})$
$=\ln{\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})}\exp(-\eta\tilde{\ell}_{t}^{i})$ (30)
$\leq\ln\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\left(1-\eta\tilde{\ell}_{t}^{i}+\frac{\eta^{2}(\tilde{\ell}_{t}^{i})^{2}}{2}\right)$ (31)
$=\ln\left(1-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})(\tilde{\ell}_{t}^{i})^{2}\right)$
$\leq-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})(\tilde{\ell}_{t}^{i})^{2}$ (32)
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\left(\frac{\ell_{t}(\tau_{i},\alpha)\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}}))}{{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}}\right)\ell_{t}(\tau_{i},\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))$ (33)
${=}-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}\frac{{\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}}))\cdot p_{t}(\tau_{i})}\cdot\ell_{t}(\tau_{i},\alpha)\cdot\ell_{t}(\tau_{i},\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$
${=}-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}})}\frac{{p_{t}(\tau_{i})}\cdot\ell_{t}(\tau_{i},\alpha)\cdot\ell_{t}(\tau_{i},\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$
${=}-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}(\tau_{t})}\frac{{p_{t}(\tau_{i})}\cdot\ell_{t}(\tau_{i},\alpha)^{2}}{(\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau}))^{2}}$
${\leq}-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}})}\frac{{p_{t}(\tau_{i})}}{(\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau}))^{2}}$
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}})}\frac{{p_{t}(\tau_{i})}}{(\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}}))\cdot p_{t}(\bar{\tau}))^{2}}$ (34)
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}})}\frac{{p_{t}(\tau_{i})}}{(\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau}))^{2}}$ (35)
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\frac{1}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}$
$=-\eta\mathbbm{E}_{\tau_{i}\sim p_{t}}\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\frac{1}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}.$
Here, (30) follows directly from the definition of the hypotheses distribution $p_{t}(\tau_{t})$, (31) is obtained by applying the second-order Taylor bound $e^{x}\leq 1+x+{x^{2}}/{2}$ which is valid for $x\leq 0$, together with the observation that $-\eta\tilde{\ell}_{t}^{i}\leq 0$, (32) hold by the standard logarithmic bound $\ln(1+x)\leq x$ for any $x\in\mathbb{R}$. (33) follows from the definition of our estimator. (34) holds as $\tau_{i},\bar{\tau}\in\mathcal{{H}}_{t}(\bar{\tau})=\mathcal{{H}}_{t}(\tau_{t})$ due to conditions in summations and (22) i.e., $\tau_{i}\in\mathcal{{H}}_{t}(\tau_{t})$ and $\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})$ in the summations derive $\mathcal{{H}}_{t}(\tau_{t})=\mathcal{{H}}_{t}(\tau_{i})=\mathcal{{H}}_{t}(\bar{\tau})$ from (22). Finally, (35) holds as $\tau_{i}\in\mathcal{{H}}_{t}(\tau_{t})$ is always true due to the outer summation condition.
Then, we have the following:
$\mathbbm{E}\left[\ln\frac{W_{T}}{W_{0}}\right]$ $=\mathbbm{E}\left[\sum_{t=1}^{T}\ln\frac{W_{t}}{W_{t-1}}\right]$
@@
$\leq\mathbbm{E}\left[-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\frac{\ell_{\text{max}}^{2}}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}\right]$ (36)
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\frac{\ell_{\text{max}}^{2}}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})},$
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+{\eta^{2}}\ell_{\text{max}}^{2}T,$ (37)
where, after the telescoping sum, (36) applies the exponential-weights bound obtained in the previous step to each factor $\ln\frac{W_{t}}{W_{t-1}}$ and (37) holds as follows:
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}$ $\frac{\ell_{\text{max}}^{2}}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}$
$=\sum_{t=1}^{T}\sum_{\tau_{t}\in\mathcal{{H}}}\frac{p_{t}(\tau_{t})\ell_{\text{max}}^{2}}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}$
$=\sum_{t=1}^{T}\ell_{\text{max}}^{2}\sum_{\tau_{t}\in\mathcal{{H}}}\frac{p_{t}(\tau_{t})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}$
@@
$=\sum_{t=1}^{T}\ell_{\text{max}}^{2}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}\right)$
@@
$=\sum_{t=1}^{T}\ell_{\text{max}}^{2}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}\right)$ (38)

### v2

@@ H.4 A Proof of Theorem 1
Also, the following equality with respect to our loss estimator (7) holds similarly:
$\mathbbm{E}_{\tau\sim p_{t}}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))$ $=\sum_{\tau\in\mathcal{{H}}}p_{t}(\tau)\frac{\ell_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}(\bar{\tau}))p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}(\tau_{t}))$
$=\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\tau)\frac{\ell_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}(\bar{\tau}))p_{t}(\bar{\tau})}$
$=\ell_{t}(\tau_{t})\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}\frac{p_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}(\bar{\tau}))p_{t}(\bar{\tau})}$ (49)
$=\ell_{t}(\tau_{t})\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}\frac{p_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\bar{\tau})}$ (50)
$=\ell_{t}(\tau_{t})\frac{\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\bar{\tau})}$
$=\ell_{t}(\tau_{t})\frac{\gamma_{t}+\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\bar{\tau})}-\ell_{t}(\tau_{t})\frac{\gamma_{t}}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\bar{\tau})}$ (51)
$=\ell_{t}(\tau_{t})-\gamma_{t}\frac{\ell_{t}(\tau_{t})}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}\mathbbm{1}(\tau_{t}\in\mathcal{{H}}_{t}(\bar{\tau}))p_{t}(\bar{\tau})}\mathbbm{1}(\tau_{t}\in\mathcal{{H}}_{t}(\tau_{t}))$ (52)
$=\ell_{t}(\tau_{t})-\gamma_{t}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t})),$ (53)
where (49) holds as $\ell_{t}(\tau)=\ell_{t}(\tau^{\prime})$ for all $\tau\in\mathcal{{H}}(\tau^{\prime})$ by the definition of the specialized loss in (4), (50) holds as the same as (47), (51) holds as $\tau_{t}\in\mathcal{{H}}_{t}(\tau_{t})$ and (43), (52) holds due to the property of the indicator function, and (53) holds by the definition.
Then, we first introduce a new Lemma 54 and Corollary 63 to establish a high probability bound.
Lemma 3.
Let $\alpha_{t}\leq 2\gamma_{t}$ where $\gamma_{t}$ is non-increasing in $t$. Then, the following inequality holds with probability at least $1-\delta$:
$\sum_{t=1}^{T}\alpha_{t}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-\sum_{t=1}^{T}{2\alpha_{t}\ell_{\text{max}}}\leq\ell_{\text{max}}\ln(1/\delta).$ (54)
Proof.
First, the following inequality holds for any $\tau\in\mathcal{{H}}$:
$\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))$ $=\frac{\ell_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))$
$=\ell_{\text{max}}\cdot\frac{\ell_{t}(\tau)/\ell_{\text{max}}}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))$
$\leq\ell_{\text{max}}\cdot\frac{\ell_{t}(\tau)/\ell_{\text{max}}}{\gamma_{t}\ell_{t}(\tau)/\ell_{\text{max}}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))$
$=\frac{\ell_{\text{max}}}{2\gamma_{t}}\cdot\frac{2\gamma_{t}\ell_{t}(\tau)/\ell_{\text{max}}}{\gamma_{t}\ell_{t}(\tau)/\ell_{\text{max}}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))$
$=\frac{1}{\beta_{t}}\cdot\frac{\frac{\beta_{t}\ell_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}}{\frac{\beta\ell_{t}(\tau)}{2\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}+1}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))$
$=\frac{1}{\beta_{t}}\cdot\frac{\frac{\beta_{t}\ell_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))}{\frac{\beta_{t}\ell_{t}(\tau)}{2\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))+1}$
$=\frac{1}{\beta_{t}}\cdot\frac{\beta_{t}\hat{\ell}_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))}{\beta_{t}\hat{\ell}_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))/2+1}$
$\leq\frac{1}{\beta_{t}}\cdot\ln\left(1+\beta_{t}\hat{\ell}_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\right),$ (55)
where $\beta_{t}=2\gamma_{t}/\ell_{\text{max}}$, $\hat{\ell}_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\coloneqq\frac{\ell_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))$, and the last inequality holds as $\frac{x}{1+x/2}\leq\ln(1+x)$ for all $x\geq 0$.
Let $\mathbbm{E}_{t}$ be the expectation conditioned on $\tau_{1},\dots,\tau_{t-1}$ and $\tau^{\prime}$ be any element in $\mathcal{{H}}$. Then, we have
$\mathbbm{E}_{t}$ $\left[\exp\left(\frac{\alpha_{t}}{\ell_{\text{max}}}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))\right)\right]$
@@
$\leq\mathbbm{E}_{t}\left[\exp\left(\frac{\alpha_{t}}{\ell_{\text{max}}\beta_{t}}\ln\left(1+\beta_{t}\hat{\ell}_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))\right)\right)\right]$ (56)
$\leq\mathbbm{E}_{t}\left[1+\frac{\alpha_{t}}{\ell_{\text{max}}}\hat{\ell}_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))\right]$ (57)
$=1+\frac{\alpha_{t}}{\ell_{\text{max}}}\mathbbm{E}_{t}\hat{\ell}_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))$
$=1+\frac{\alpha_{t}}{\ell_{\text{max}}}\sum_{\tau_{t}\in\mathcal{{H}}}p_{t}(\tau_{t})\frac{\ell_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))$
$=1+\frac{\alpha_{t}}{\ell_{\text{max}}}\sum_{\tau_{t}\in\mathcal{{H}}}p_{t}(\tau_{t})\frac{\ell_{t}(\tau_{t})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\bar{\tau})}$
$=1+\frac{\alpha_{t}}{\ell_{\text{max}}}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau^{\prime})}\frac{p_{t}(\tau)\ell_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau^{\prime})}\frac{p_{t}(\tau)\ell_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}\right)$
$\leq 1+\frac{\alpha_{t}}{\ell_{\text{max}}}\ell_{\text{max}}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau^{\prime})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau^{\prime})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}\right)$
$=1+\alpha_{t}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau^{\prime})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau^{\prime}})}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau^{\prime})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}\right)$ (58)
$=1+\alpha_{t}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau^{\prime})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau^{\prime}})}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau^{\prime})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in{\mathcal{{H}}_{t}^{C}(\tau^{\prime})}}p_{t}(\bar{\tau})}\right)$ (59)
$\leq 1+2\alpha_{t}$ (60)
$\leq\exp(2\alpha_{t}),$ (61)
where (56 holds as (55). (57) uses $x\ln(1+y)\leq\ln(1+xy)$ for all y>1y>-1 and $x\in[0,1]$ since $\alpha_{t}\leq 2\gamma_{t}=\ell_{\text{max}}\beta_{t}$. (58) holds as (41), i.e., $\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}(\tau_{t})$ due to $\tau\in\mathcal{{H}}_{t}(\tau_{t})$ in the summation. Also, (59) holds due to (42), i.e., $\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}^{C}(\tau_{t})$ as $\tau\in\mathcal{{H}}_{t}^{C}(\tau_{t})$ from the summation condition. Finally, (60) holds with the equality unless $\mathcal{{H}}_{t}(\tau)=\emptyset$ or $\mathcal{{H}}_{t}^{C}(\tau)=\emptyset$, i.e., $f_{t}=0$, in which case the inequality holds.
Thus, we can find that $W_{t}\coloneqq\exp\left(\sum_{s=1}^{t}\frac{\alpha_{s}}{\ell_{\text{max}}}\left(\ell_{s}(\tau_{s}\mid\mathcal{{H}}_{s}(\tau_{s}))-2\ell_{\text{max}}\right)\right)$ is a supermartingale with respect to $\mathbbm{E}_{t}$, i.e., $\mathbbm{E}_{t}\left[W_{t}\right]\leq W_{t-1}$ as follows:
$\mathbbm{E}_{t}[W_{t}]$ $=\mathbbm{E}_{t}\left[W_{t-1}\exp\left(\frac{\alpha_{s}}{\ell_{\text{max}}}\left(\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-2\ell_{\text{max}}\right)\right)\right]$
$=W_{t-1}\mathbbm{E}_{t}\left[\exp\left(\frac{\alpha_{s}}{\ell_{\text{max}}}\left(\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-2\ell_{\text{max}}\right)\right)\right]$
$=W_{t-1}\mathbbm{E}_{t}\left[\exp\left(\frac{\alpha_{s}}{\ell_{\text{max}}}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))\right)\right]\exp\left(-2\alpha_{s}\right)$
$\leq W_{t-1}\exp\left(2\alpha_{t}\right)\exp\left(-2\alpha_{s}\right)$ (62)
$=W_{t-1},$
where (62) holds due to (61). Due to $W_{0}=1$ by the definition, the above implies $\mathbbm{E}\left[W_{T}\right]\leq 1$. Then, by applying this with the Markov’s inequality, we have
@@
$\mathbbm{P}\left(\sum_{t=1}^{T}\frac{\alpha_{t}}{\ell_{\text{max}}}\left(\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-2\ell_{\text{max}}\right)>\ln\frac{1}{\delta}\right)=\mathbbm{P}\left(W_{T}>\frac{1}{\delta}\right)\leq\delta\mathbbm{E}_{T}[W_{T}]\leq\delta.$
Thus,
$\sum_{t=1}^{T}\frac{\alpha_{t}}{\ell_{\text{max}}}\left(\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-2\ell_{\text{max}}\right)\leq\ln(1/\delta),$
with probability at least $1-\delta$, which completes the proof. ∎
Corollary 2.
Let $\gamma_{t}=\gamma>0$ for all $t$. Then, simultaneously for all $\tau\in\mathcal{{H}}$,
$\sum_{t=1}^{T}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))-\sum_{t=1}^{T}\ell_{t}(\tau)\leq\frac{\ell_{\text{max}}\ln(|\mathcal{{H}}|/\delta)}{2\gamma},$ (63)
with probability at least $1-\delta$.
Proof.
The proof of this corollary can be derived from Lemma 54, but here we provide a direct proof, which uses similar proof techniques as in Lemma 54.
Let $\gamma_{t}=\gamma\geq 0$. Then, we have
@@
$\mathbbm{E}_{t}\left[\exp\left(\frac{2\gamma}{\ell_{\text{max}}}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\right)\right]$ $\leq\mathbbm{E}_{t}\left[\exp\left(\frac{2\gamma}{\ell_{\text{max}}\beta}\ln\left(1+\beta\hat{\ell}_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\right)\right)\right]$ (64)
$\leq\mathbbm{E}_{t}\left[1+\frac{2\gamma}{\ell_{\text{max}}}\hat{\ell}_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\right]$ (65)
$=1+\frac{2\gamma}{\ell_{\text{max}}}\ell_{t}(\tau)$ (66)
$\leq\exp\left(\frac{2\gamma}{\ell_{\text{max}}}\ell_{t}(\tau)\right),$ (67)
where (64) holds same as (55) with fixed $\gamma_{t}=\gamma$ and $\beta_{t}=\beta$ for all $t$. Also, (65) uses $x\ln(1+y)\leq\ln(1+xy)$ for all y>1y>-1 and $x\in[0,1]$ along with the fact that $2\gamma=\ell_{\text{max}}\beta$ and thus $\frac{2\gamma}{\beta\ell_{\text{max}}}=1$. (66) holds as $\mathbbm{E}_{t}\hat{\ell}_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))=\ell_{t}(\tau)$ from 48.
Thus, we can find that $W_{t}\coloneqq\exp\left(\sum_{s=1}^{t}\frac{2\gamma}{\ell_{\text{max}}}\left(\ell_{s}(\tau\mid\mathcal{{H}}_{s}(\tau_{s}))-\ell_{s}(\tau)\right)\right)$ is a supermartingale with respect to $\mathbbm{E}_{t}$, i.e., $\mathbbm{E}\left[W_{t}\right]\leq W_{t-1}$ as follows: